OpenStudy (anonymous):

<17 and <18 are vertical angles. m <17=(5x+3)(degrees) and m<18=(3x+23)(degrees) find the measure of <18 a. 13 degrees b. 44.3 degrees c. 53 degrees d. 33 degrees. i just need the answer please i have to go!

OpenStudy (anonymous):

angle 17 and angle 18 are equal to each other for vertical angles are equal :P so 5x + 3 = 3x + 23 can you combine like terms?

OpenStudy (anonymous):

yes 5x and 3x?

OpenStudy (anonymous):

no o.o 5x + 3 = 3x + 23 - 3 - 3 5x = 3x + 20 -3x -3x 2x = 20 2x/2 = 20/2 x = 11

OpenStudy (anonymous):

is it 53??